We can write the demand and supply curves algebraically as follows:
Demand: Q = a – bP
Supply: Q = c + dP
a) Find the values of the constants a, b, c, and d in the supply and demand equations based
on the supply and demand for the world copper market. The relevant numbers for the
copper market are as follows: Equilibrium Quantity Q∗= 18 million metric tons per year
(mmt/yr) Equilibrium Price P∗ = $3.00 per pound, Elasticity of supply ES = 1.5 Elasticity
of demand ED = -0.5.
1
Expert's answer
2021-01-05T07:54:44-0500
"a-3b=18"


"E_D=-b\frac{p}{Q}"


"b=3"


"a=27"

"Q_D=27-3p"


"c+3d=18"


"E_S=d\frac{3}{18}"


"d=9"


"c=-9"


"Q_S=-9+9p"